In 2000, a company had 1148 stores nationwide. By 2002, this total had grown to 1542. If the number of stores continues to grow exponentially at the same rate, how many stores will be there in 2016?
Answer by jorel1380(3719)

P(2002)=P(2000)*e^rt
1542=1148*e^rt=1148*e^2r
r=0.14752948862188153567884565687251
So:
P(16)=1148*e^16*0.14752948862188153567884565687251
P(2016)=12164 stores
☺☺☺☺